# XARP: A Human-First and AI Agent-Ready Extended Reality Toolkit

### v2

@@ 5.1 Participants
5.2 Measures
We assessed the potential acceptance of XARP using the Unified Theory of Acceptance and Use of Technology (UTAUT) scale [48], a widely used instrument with strong psychometric validity for evaluating behavioral intention to adopt a technology. Because the Social Influence construct primarily applies in organizational contexts with mandatory use, and XARP adoption is expected to be voluntary, we excluded this construct. We kept the remaining constructs: Performance Expectancy, Effort Expectancy, Facilitating Conditions, and the outcome Behavioral Intention. In addition, we collected open-ended responses asking participants to explain their reasons for adopting or avoiding XARP.
To assess the perceived value of the toolkit, we designed a custom survey based on the HCI Toolkit Goals (HCITG) compiled by Ledo et al. [28]. Since no standardized survey instrument exists to assess the toolkit goals proposed in this framework, we constructed a custom instrument named HCITG. We mapped each toolkit goal compiled by Ledo et al. into a survey construct, further separating creative exploration from solution replication. Each HCITG survey construct was measured with three to six Likert-scale items, including at least one reverse-coded item. In summary, our customized HCITG instrument included: Reducing Authoring Time, Creating Paths of Least Resistance, Empowering New Audiences, Integrating with Current Practices and Infrastructure.
@@ 5.4 Data Analysis
We analyzed participant answers about reasons to adopt or avoid XARP, and their open-ended questions, using a rapid variant of the Framework Method [19]. This approach provides a transparent, systematic way to code and chart even small corpora, enabling clear comparison across responses while minimizing analytic overhead.
6 Results
6.1 Potential Acceptance
@@ 6.3 Toolkit Comprehension
A rapid Framework analysis [19] of participant responses to the question, “What other questions do you have that were not answered in the videos?” produced seven main FAQ topics. The most frequent topic was Compatibility & Integration (42%), e.g., “Is XARP compatible with other AR headsets beyond the Quest?” and “Does the [toolkit] have anything for integrating with Unity Sentis and on-device AI models?” Another frequent topic was Limitations (15%), e.g., “(…) In what circumstances have you found the built-in commands, LLM prompting, or motion/data tracking to be unreliable?”
7 Discussion
7.1 Factors of AI-XR Toolkit Adoption
7.1.1 Performance Expectancy and Authoring Time Reduction
@@
7.1.2 Facilitating Conditions and Paths of Least Resistance
@@
7.2 Threshold and Ceiling at Different Expertise Levels
Threshold, the effort required to begin productive use of a system, and ceiling, the sophistication of what can be achieved with the system, are guiding themes in toolkit research [38]. To examine how users perceive XARP along these dimensions, we analyzed responses from participants at two levels of expertise. Less experienced participants may emphasize challenges associated with the toolkit threshold, while those with greater expertise can offer critical feedback based on advanced use cases that delimit the toolkit ceiling.
@@ 7.2 Threshold and Ceiling at Different Expertise Levels
Experts rated all constructs in HCITG positively, and there was a greater variance in ratings for Creative Exploration compared to the other constructs. This is consistent with participant comments that highlighted the hyper-specific use cases that most experts engage with, and suggests varied interpretations of “creative exploration” across the experts in this study. For example, P12 valued the “Ease of prototyping, [allowing me to] benchmark a bunch of API callable models quickly on a specific task.”, and P23 liked “The ability to iterate quickly due to the server-first nature”. Other experts found XARP limiting in it’s ability to support features such as “visual design and tailored immersive experiences” (P19), raised concerns about limitations due to API token usage (P5, P7) and a general lack of low-level flexibility (P20). This reflects the tradeoff between convenient abstractions that support quick development and fine-grained control [46].
7.3 A Future Vision for the AI-XR Ecosystem
Unlike previous strategies to address complexity in software, developers can now delegate design decisions to AI agents, essentially trading control over specification for adaptive behavior at runtime. This capability enables a new scaling mechanism, unconstrained by the burden of upfront human design. Acting as personal co-designers, agents can generate bespoke solutions at runtime. If made efficient, such agents could be deployed directly to end-users, generating unique applications on demand. In this vision, developers do not vanish, they climb the “abstraction ladder” [25]. Their role shifts from delivering finished products to shaping human–AI co-design workflows, embedding human insight into hybrid design teams, and consolidating patterns that emerge from millions of personalized solutions. Business logic will change accordingly. Rather than competing through standardized products, companies could offer design-as-a-service, deploying AI agents with distinct design styles. A marketplace of such agents may emerge, with interoperability standards like MCP enabling both integration and economic exchange.
@@ 7.3 A Future Vision for the AI-XR Ecosystem
Our vision builds on a long-standing effort to empower end-users. Alan Kay’s proposal of the “personal dynamic medium” set the precedent for imagining computing as something users could directly shape rather than merely consume [27]. By the 1990s, the HCI community had begun exploring end-user programming as a path toward broadening participation in software creation [39]. In the 2000s, this agenda matured into end-user software engineering, articulating the technical and methodological challenges of enabling non-professionals to construct reliable software [9]. AI agents extend this trajectory, potentially democratizing software creation even further.
8 Limitations & Future Work
Our design pattern analysis aimed for representativeness, but our search may have excluded relevant papers. While our walkthrough study recruitment process engaged participants with varied expertise across five countries, the reliance on personal distribution channels introduces potential sampling bias, and the sample size may have constrained statistical power. The walkthrough study was suitable for assessing early acceptance and perceived value of XARP, but further research should evaluate direct use of more mature versions in hybrid workshops and take-home studies [28].

### v3

@@ 6.1. Participants
6.2. Measures
We assessed the potential acceptance of XARP using the Unified Theory of Acceptance and Use of Technology (UTAUT) scale (Venkatesh et al., 2003), a widely used instrument with strong psychometric validity for evaluating behavioral intention to adopt a technology. Because the Social Influence construct primarily applies in organizational contexts with mandatory use, and XARP adoption is expected to be voluntary, we excluded this construct. We kept the remaining constructs: Performance Expectancy, Effort Expectancy, Facilitating Conditions, and the outcome Behavioral Intention. In addition, we collected open-ended responses asking participants to explain their reasons for adopting or avoiding XARP.
To assess the perceived value of the toolkit, we designed a custom survey based on the HCI Toolkit Goals (HCITG) compiled by Ledo et al. (Ledo et al., 2018). Since no standardized survey instrument exists to assess the toolkit goals proposed in this framework, we constructed a custom instrument named HCITG. We mapped each toolkit goal compiled by Ledo et al. into a survey construct, further separating creative exploration from solution replication. Each HCITG survey construct was measured with three to six Likert-scale items, including at least one reverse-coded item. In summary, our customized HCITG instrument included: Reducing Authoring Time, Creating Paths of Least Resistance, Empowering New Audiences, Integrating with Current Practices and Infrastructure.
@@ 6.4. Data Analysis
We analyzed participant answers about reasons to adopt or avoid XARP, and their open-ended questions, using a rapid variant of the Framework Method (Gale et al., 2013). This approach provides a transparent, systematic way to code and chart even small corpora, enabling clear comparison across responses while minimizing analytic overhead.
7. Results
7.1. Potential Acceptance
@@ 7.3. Toolkit Comprehension
A rapid Framework analysis (Gale et al., 2013) of participant responses to the question, “What other questions do you have that were not answered in the videos?” produced seven main FAQ topics. The most frequent topic was Compatibility & Integration (42%), e.g., “Is XARP compatible with other AR headsets beyond the Quest?” and “Does the [toolkit] have anything for integrating with Unity Sentis and on-device AI models?” Another frequent topic was Limitations (15%), e.g., “(…) In what circumstances have you found the built-in commands, LLM prompting, or motion/data tracking to be unreliable?”
8. Discussion
8.1. Factors of AI-XR Toolkit Adoption
8.1.1. Performance Expectancy and Authoring Time Reduction
@@
8.1.2. Facilitating Conditions and Paths of Least Resistance
@@
8.2. Threshold and Ceiling at Different Expertise Levels
Threshold, the effort required to begin productive use of a system, and ceiling, the sophistication of what can be achieved with the system, are guiding themes in toolkit research (Myers et al., 2000). To examine how users perceive XARP along these dimensions, we analyzed responses from participants at two levels of expertise. Less experienced participants may emphasize challenges associated with the toolkit threshold, while those with greater expertise can offer critical feedback based on advanced use cases that delimit the toolkit ceiling.
@@ 8.2. Threshold and Ceiling at Different Expertise Levels
Experts rated all constructs in HCITG positively, and there was a greater variance in ratings for Creative Exploration compared to the other constructs. This is consistent with participant comments that highlighted the hyper-specific use cases that most experts engage with, and suggests varied interpretations of “creative exploration” across the experts in this study. For example, P12 valued the “Ease of prototyping, [allowing me to] benchmark a bunch of API callable models quickly on a specific task.”, and P23 liked “The ability to iterate quickly due to the server-first nature”. Other experts found XARP limiting in it’s ability to support features such as “visual design and tailored immersive experiences” (P19), raised concerns about limitations due to API token usage (P5, P7) and a general lack of low-level flexibility (P20). This reflects the tradeoff between convenient abstractions that support quick development and fine-grained control (Spolsky, 2004).
8.3. A Future Vision for the AI-XR Ecosystem
Unlike previous strategies to address complexity in software, developers can now delegate design decisions to AI agents, essentially trading control over specification for adaptive behavior at runtime. This capability enables a new scaling mechanism, unconstrained by the burden of upfront human design. Acting as personal co-designers, agents can generate bespoke solutions at runtime. If made efficient, such agents could be deployed directly to end-users, generating unique applications on demand. In this vision, developers do not vanish, they climb the “abstraction ladder” (Jr., 1987). Their role shifts from delivering finished products to shaping human–AI co-design workflows, embedding human insight into hybrid design teams, and consolidating patterns that emerge from millions of personalized solutions. Business logic will change accordingly. Rather than competing through standardized products, companies could offer design-as-a-service, deploying AI agents with distinct design styles. A marketplace of such agents may emerge, with interoperability standards like MCP enabling both integration and economic exchange.
@@ 8.3. A Future Vision for the AI-XR Ecosystem
Our vision builds on a long-standing effort to empower end-users. Alan Kay’s proposal of the “personal dynamic medium” set the precedent for imagining computing as something users could directly shape rather than merely consume (Kay, 1977). By the 1990s, the HCI community had begun exploring end-user programming as a path toward broadening participation in software creation (Nardi, 1993). In the 2000s, this agenda matured into end-user software engineering, articulating the technical and methodological challenges of enabling non-professionals to construct reliable software (Burnett et al., 2004). AI agents extend this trajectory, potentially democratizing software creation even further.
9. Limitations & Future Work
Our design pattern analysis aimed for representativeness, but our search may have excluded relevant papers. While our walkthrough study recruitment process engaged participants with varied expertise across five countries, the reliance on personal distribution channels introduces potential sampling bias, and the sample size may have constrained statistical power. The walkthrough study was suitable for assessing early acceptance and perceived value of XARP, but further research should evaluate direct use of more mature versions in hybrid workshops and take-home studies (Ledo et al., 2018).